# Pulsed magnetic field gradient on a tip for nanoscale imaging of spins

### v1

@@ Abstract
Nanoscale magnetic resonance imaging (nanoMRI) aims at obtaining structure at the single molecule level. Most of the techniques for effecting a nanoMRI gradient use small permanent magnets. Here, we present a switchable magnetic field gradient on a tip, which is designed to provide a local and controllable magnetic field with a high gradient on the nanometer scale. We incorporate the gradient field with a nanoscale magnetic resonance sensor, a single nitrogen-vacancy (NV) center in diamond, to provide high-resolution magnetic resonance imaging. The device is a metal microwire deposited along a quartz tip, with the current flowing along the tip inducing a magnetic field around its apex. This field can be manipulated throughout a measurement by controlling the current along the wire. We achieved gradients as high as 1 at fields weaker than 200 . Such a gradient can facilitate electron spin mapping with 1 nm resolution using single NV sensors, allowing for nanoscale imaging of electrons. The ability to switch the current on and off and to position the device with high precision overcomes limitations such as limited emitter contrast and the flexibility in sample preparation. Moreover, we show that proximity of the metallic tip to the sensor modifies the Rabi power in a spatially dependent manner, providing regions with enhanced (3.5) and decreased Rabi power. This spatial gradient, induced by the tip, offers the opportunity for selective pulses on nearby spin species where the same microwave power will result in different spin manipulation characteristics.
keywords:
AFM, nanoMRI, NV center in diamond, Rabi power, quantum sensing1 Introduction
@@
Nanoscale magnetic resonance imaging (nanoMRI) aims to give insight to molecular structure at the single molecule level, and has advanced significantly over the past decades [1]. It has the potential to gain deeper understanding in the vast worlds of chemistry and biology where conventional nuclear magnetic resonance (NMR) is limited by sample size [2] and conventional MRI is limited by magnetic field gradients to the micron scale [3]. The basic principle of MRI is the encoding of spatial information via spectral variation of the spin’s Larmor precession frequency. In this way, a known gradient magnetic field is applied over the detection region. This separates the resonance frequency of similar species in the sample via , where is the gyromagnetic ratio of the detected species and is the local magnetic field magnitude. A variety of techniques exist in the field of nanoMRI, among them - magnetic resonance force microscopy (MRFM), electron spin resonance using scanning tunneling microscopy (ESR-STM) and the nitrogen-vacancy (NV) center in diamond. Modern MRFM cantilevers allow detection of on the order of molecules [1] with an impressive 10 nm resolution [4, 5]. Moreover, and specifically relevant to our work, a recent method of applying the gradient was implemented by focusing the current in a planar device below the cantilever, resulting in resolution below 2 nm [6]. ESR-STM is a surface technique and is limited by the type of systems that can currently be explored [7]. Both MRFM and ESR-STM are technically very challenging.
@@
In NV magnetic resonance sensing, the spin state of the NV interacts with magnetic fields in its immediate environment and the field is measured via optical detection of the NV spin state [8]. Nanoscale resolution can be achieved by sensing single NV defects which can detect local magnetic fields and have demonstrated sensitivity of [9], comparable to sensing a single electron magnetic field of at a distance of 10 nm. Further experiments have detected NMR signals with nanoscale sample size detecting as few as nuclear spins [10, 11], and later on reaching the single nuclear spin limit [12, 13]. In these experiments, the signal is integrated over all spins interacting with the NV, and one is unable to differentiate between individual spins. The gradient-based spatial resolution, as in MRI, was demonstrated by Grinolds et al. [14] where sub-nm resolution was reached by depositing a thin ferromagnetic film on a tip which applied a local gradient magnetic field [15]. The spatial gradient gives rise to spectral separation that encodes the location of nearby molecules within the NV sensing region. This impressive milestone is limited by the strong off-axis magnetic fields induced by the ferromagnet, as such fields are detrimental to NV contrast due to mixing of the zero-field splitting (ZFS) eigenstates. The ZFS defines the NV’s -axis [16]. This can be seen in the ground-state spin Hamiltonian of the NV center,
@@
with being the ground-state ZFS, the magnetic field along the NV axis, is the Bohr magneton and is the Landé g-factor for the NV. Specifically, it is these off-axis () fields that are created by a ferromagnet and are undesired during the initialization and readout stages of the NV’s pulse sequence used for sensing.
@@
Current-based pulsed magnetic fields can overcome this challenge, as they can be switched on and off throughout the experiment. Thus, the magnetic field is applied during the sensing stage of the pulse sequence but switched off during spin read-out, which is optimal when the NV eigenstates are prominently defined by ZFS and terms. Switching the current using an electromagnet was demonstrated by Bodenstedt et al. [17] where a hard drive’s write head was used to achieve magnetic gradients up to 100 , yet also here magnetic noise from the write head’s electromagnet limited the detection contrast. An additional advantage of pulsed magnetic fields is that they allow one to measure small shifts that may be masked by drifts in the system. With a switchable field, one can sequentially measure a signal with and without the magnetic field, thus becoming indifferent to system drifts on longer time scales. Lastly, this allows unique pulse sequences that leverage a changing gradient within an experiment, such as Fourier imaging [18, 19, 20] and gradient pulse sequences [21, 22]. Therefore, a switchable magnetic field gradient is a requisite for atomic-scale NV-MRI. In this work, we demonstrate a device tailored for nanoMRI, that provides an on-demand localized magnetic field gradient (see Fig. 1a).
2 Current focusing device design
@@
Inspired by the MRFM community [6], the device we present was developed in order to achieve controllable magnetic field gradients localized around a single NV, which acts as a magnetic sensor. To this end, we fabricate a wire by physical vapor deposition (see Methods) that runs along a quartz tip. The abrupt change in the wire’s angle at the apex of the tip results in a gradient in the magnetic field induced by the current running through the wire. The field of such a configuration is estimated by numerical integration according to the Biot-Savart law (see SI).
@@
The simulation results are presented in Fig. 1(b,c), showing the magnitude of the magnetic field in-plane and out-of-plane relative to the tip. The field strength is weak relative to the zero field splitting, and so this field is not expected to impact the eigenstates of the NV [16]. We considered a working point of 50 nm as a stand-off distance between the tip and the diamond, where the maximum field strength is 0.89 mT (see SI). The maximum gradient will determine the spatial resolution, calculated to be when running a current of 2 mA. The final resolution of the system depends also on the spectral resolution, or linewidth, of the sensor, for a typical single NV setup at ambient conditions this is around 100 kHz for shallow implantation, corresponding to a dephasing time of of , where ) [23]. Hence, the expected spatial resolution for electron spins (with the electron’s gyromagnetic ratio taken to be GHz/T) is .
@@
We found the maximal current in the device is limited due to Joule heating induced by the current, which can cause device failure, i.e., open circuit. In order to estimate the heating of the device, we simulate the heating using COMSOL to solve the heat transfer equations for the device geometry (see SI). From the perspective of heating, a larger cross-section is beneficial to sustain increasing currents (see Fig. LABEL:fig:temperature), thus a larger apex will prevent device failure due to heating. Nevertheless, our simulations (as shown in Fig. 1) suggest that a smaller apex diameter leads to a stronger gradient. The chosen geometry, therefore, of the device, was set to balance the need for low heating while still achieving a gradient that is beneficial for magnetic resonance imaging. To this end, the tip is designed to have an apex on the order of 1 m and a deposition thickness on the order of 100 nm (see Methods).
3 The gradient sensing setup
@@
4 Magnetic gradient characterization
@@
To map the DC magnetic field of the tip, an optically detected magnetic resonance (ODMR) pulse sequence, sensitive to magnetic fields that are projected along the NV zero-field splitting axis, is used, where current is applied during the spin state manipulation but turned off during the spin readout and initialization phase, see Fig. 2c. We identify a rise/fall time of 600 ns (see SI for the MW/tip circuitry). The pulse sequence is alternated with and without current to cancel the effect of slow drifts that would otherwise accumulate and mask small changes in the magnetic field. The tip is scanned relative to the NV and ODMR is measured at each tip location. To avoid damage to the tip, the measurements were performed with a current of 1.54 mA, as measured on the oscilloscope (see SI for a discussion on the current amplitude).
@@
The magnetic field measured at a single pixel (tip position) is shown in Fig. 3, with each resonant dip fit to a Gaussian. The shift in the magnetic field due to the tip is calculated as the difference between the center dip measured with and without the current. The magnetic field is then measured over a lateral range of around the pillar with an NV. Scans performed at varying distances to the diamond’s surface are shown in Fig. 3(a,b). To compare to the simulation of the expected field (as shown in Fig. 1c), we account for the degrees of freedom in the measurement. The first is the relative angle between the NV axis to the direction of the current, and the other being the lab coordinates along which the tip is scanned. The magnitude and range of the measured field are affected by the angle between the current to the NV, since the ODMR measurement is most sensitive to fields parallel to the NV axis. While the scanning axis defines the axis connecting the maxima and minima, it does not affect the field strength. The simulation, as shown in Fig. 3, projected along the NV axis and rotated by 45∘, is qualitatively similar to the measurement. Several features are visible in both - a sharp peak in the field along with a broad yet shallow dip. As expected, the field strength increases with the distance to the surface. Quantitatively, the range and values of the field in the experiment and the simulation are in good agreement. The discrepancies can be associated with several uncertainties - the discreteness of the measurement, the error in the distance to the surface and simulation assumptions regarding the current flow along the wire. From the comparison, we can then calculate the magnetic field induced by the tip at the sample’s position. This is important when considering the effect of the magnetic field on spins external to the diamond, such as unpaired electrons on the surface and more broadly a sample placed on the diamond, since by knowing the field generated by the gradient at each point, we can deduce what the Larmor frequency of the target spin would be.
@@
To measure the magnetic gradient at higher resolution, we measure ODMR along a line connected the minima to maxima regions, see Figs. 3c and 4a. This scan was repeated as a function of distance to the diamond surface, as expected the gradient increases as the tip is closer to the diamond surface, see Fig. 4(d). For a current of 1.54 mA through the tip, the maximum measured gradient is , measured when scanning the tip along the axis connecting the minima to maxima regions. The limit on resolution is a combination of the spectral resolution of the setup and the gradient achieved by the tip. With the single NV’s dephasing time () of , we were able to measure ODMR within an error of 100 kHz. The signal from an electron is detected by the NV at the electron resonance frequency at , where and is the magnetic field applied to the electron. Thus, to distinguish electrons in our system, the magnetic field difference required is . Protocols for NV sensing with a resolution as good as have been demonstrated [25, 26], and applying such a sequence in combination with a gradient field can improve the spatial resolution further. Finally, while we performed most of the measurements at a relatively low current, the device can support higher current densities, with a 5-fold improvement in the gradient (see SI), making it practical for nanoMRI. In fact, our low-current gradient can already distinguish between spin labels, for example, spaced 4.8 nm in a molecular ruler [27], with a Larmor precession frequency difference between them of 134 kHz.
5 Tip-induced Rabi power modulation
@@
Interestingly, the geometry used here, i.e., where a metallic tip is placed above the diamond while a metallic microwave waveguide is placed below it, is shown to affect the NV Rabi period dramatically. We observe that by a proper positioning of the metallic tip, the Rabi oscillations’ frequency (, the Rabi driving power) can be increased for a set MW power applied to the waveguide. Fig. 5 shows two sets of measurements demonstrating this effect. First, we scan the distance to the diamond surface for a set lateral position and show a relative, 3.5-fold increase in Rabi power (Fig. 5a). Next, Fig. 5b shows a lateral scan where the tip is set to a distance of 156 nm from the surface, and we plot the ratio between the measured Rabi power, , and the Rabi power when the tip is far-away from the NV, . Here, there are three noticeable regions - the far-field where the Rabi power is unaffected and equal to the baseline Rabi power as measured without a tip; an enhanced region where the Rabi power increases relative to the baseline, and; a third region where the Rabi power is significantly weaker relative to the baseline.
@@
The observed modulation of the Rabi power is due to the metallic tip modulating the magnetostatic near-fields that interact with the NV center. Since these are near-field interactions, the modulation’s strength is very sensitive to the tip’s position. A detailed analytical model, representing the tip geometry using oblate ellipsoids is given in the SI. The contours in Fig. 5b are iso-power lines calculated according to the model, and are in agreement with the experimental data. Our model also qualitatively captures the vertical distance dependence (orange line in Fig. 5a). When simplifying each ellipsoid to a single point dipole (see SI), the data can be fit by a behavior (green curve in Fig. 5a), where is the tip’s vertical distance from the NV and is the lateral one.
@@
6 Conclusions
We introduced a magnetic-field gradient on an AFM tip which can be pulsed with rise/fall times shorter than 600 ns. This is achieved using a three-step self-aligning deposition process, effectively creating a focused current at the apex of the tip. Using an NV center as a local magnetic sensor, we demonstrated gradients as strong as , allowing for nanoscale magnetic resonance imaging of proximate unpaired electrons spins in, e.g., a molecule. Moreover, this metallic tip can also serve as a means to locally modulate the driving field around both the target and sensor (NV) spin, yielding an enhancement of in the spin’s Rabi driving power. The spatial variability of the Rabi power is and of itself a gradient source for future magnetic resonance imaging in nanometer length scales. The current focusing device presented here offers several advantages. First, the ability to pulse the current allows for different magnetic fields during spin manipulation, readout, and initialization, providing optimal NV contrast as defined by the diamond lattice eigenstates. Second, the magnetic field induced by the tip is small relative to the external field and the zero field splitting, meaning only minor shifts are expected, and so the effective Zeeman energy remains largely unchanged when switching between measurements with and without the field from the tip. Lastly, the current source can also reverse the direction of the magnetic field, enabling the application of gradient pulse sequences similar to those used in conventional NMR [21, 22]. The possibility to co-align the tip with an arbitrary NV also holds advantages to sensing when working with single NVs. With the geometry presented here, a sample can be placed on the diamond surface in a statistical manner and once an optimal NV is identified, the gradient can be applied locally around the specified NV. Combined with optimal thermal management, optimal apex shape and additional protocols for nanoscale MRI, the demonstrated technology allows for an on-demand application of a magnetic resonance gradient, and is readily-compatible with current schemes for structure determination on the nanometer scale.

### v2

@@ Abstract
Nanoscale magnetic resonance imaging (nanoMRI) is crucial for advancing molecular-level structural analysis, yet existing techniques relying on permanent magnets face limitations in controllability and resolution. This study addresses the gap by introducing a switchable magnetic field gradient on a scanning tip, enabling localized, high-gradient magnetic fields at the nanoscale. Here, we demonstrate a device combining a metal microwire on a quartz tip with a nitrogen-vacancy (NV) center in diamond, achieving gradients up to 1 at fields below 200 T. This allows electron spin mapping with 1 nm resolution, overcoming challenges like emitter contrast and sample preparation rigidity. The current-controlled gradient, switchable in 600 ns, enhances precision and flexibility. Additionally, the metallic tip modifies Rabi power spatially, enabling selective spin manipulation with varying microwave effects. This innovation paves the way for advanced nanoMRI applications, including high-resolution imaging and targeted spin control in quantum sensing and molecular studies.
1 Introduction
@@
Nanoscale magnetic resonance imaging (nanoMRI) aims to give insight to molecular structure at the single molecule level, and has advanced significantly over the past decades [1]. It has the potential to gain deeper understanding in the vast worlds of chemistry and biology where conventional nuclear magnetic resonance (NMR) is limited by sample size [2] and conventional MRI is limited by magnetic field gradients to the micron scale [3]. The basic principle of MRI is the encoding of spatial information via spectral variation of the spin’s Larmor precession frequency. In this way, a known gradient magnetic field is applied over the detection region. This separates the resonance frequency of similar species in the sample via , where is the gyromagnetic ratio of the detected species and is the local magnetic field magnitude. A variety of techniques exist in the field of nanoMRI, among them - magnetic resonance force microscopy (MRFM), electron spin resonance using scanning tunneling microscopy (ESR-STM) and the nitrogen-vacancy (NV) center in diamond. Modern MRFM cantilevers allow detection of on the order of molecules [1] with an impressive 10 nm resolution [4, 5]. Moreover, and specifically relevant to our work, a recent method of applying the gradient was implemented by focusing the current in a planar device below the cantilever, resulting in resolution below 2 nm [6]. ESR-STM is a surface technique and is limited by the type of systems that can currently be explored [7]. Together with ESR-AFM [8], MRFM and ESR-STM are all technically very challenging.
@@
In NV magnetic resonance sensing, the spin state of the NV interacts with magnetic fields in its immediate environment and the field is measured via optical detection of the NV spin state [9]. Nanoscale resolution can be achieved by sensing single NV defects which can detect local magnetic fields and have demonstrated AC sensitivity better than 10 nT Hz-1/2 [10], comparable to sensing the magnetic field of a single electron at a distance of 50 nm [11]. Further experiments have detected NMR signals with nanoscale sample size detecting as few as nuclear spins [12, 13], and later on reaching the single nuclear spin limit [14, 15]. In these experiments, the signal is integrated over all spins interacting with the NV, and one is unable to differentiate between individual spins. One can use the gradient created by the NV center’s own dipole moment, thereby modifying the Larmor precession frequency of nearby spins due to the hyperfine interaction [16, 17, 18]. However, switching the gradient on and off in this case depends on the NV’s spin state and in addition the gradient itself cannot be moved around. Another gradient-based method with high spatial resolution, similar to MRI, was demonstrated by Grinolds et al. [19], where sub-nm resolution was achieved by depositing a thin ferromagnetic film on a tip which applied a local gradient magnetic field [20]. The spatial gradient gives rise to spectral separation that encodes the location of nearby molecules within the NV sensing region. This impressive milestone is limited by the strong off-axis magnetic fields induced by the ferromagnet, as such fields are detrimental to NV contrast due to mixing of the zero-field splitting (ZFS) eigenstates. The ZFS defines the NV’s -axis [21]. This can be seen in the ground-state spin Hamiltonian of the NV center,
@@
with being the ground-state ZFS, the magnetic field along the NV axis, is the Bohr magneton and is the Landé g-factor for the NV. Specifically, it is these off-axis () fields that are created by a ferromagnet that are detrimental to the initialization and readout stages of the NV’s pulse sequence used for sensing.
@@
Current-based pulsed magnetic fields can overcome this challenge, as they can be switched on and off throughout the experiment. Thus, the magnetic field is applied during the sensing stage of the pulse sequence but switched off during spin read-out, which is optimal when the NV eigenstates are prominently defined by ZFS and terms. Switching the current using an electromagnet was demonstrated by Bodenstedt et al. [22] where a hard drive’s write head was used to achieve magnetic gradients up to 100 , yet also here magnetic noise from the write head’s electromagnet limited the detection contrast. An additional advantage of pulsed magnetic fields is that they allow one to measure small shifts that may be masked by drifts in the system. With a switchable field, one can sequentially measure a signal with and without the magnetic field, thus becoming indifferent to system drifts on longer time scales. Lastly, this allows unique pulse sequences that leverage a changing gradient within an experiment, such as Fourier imaging [23, 24, 25] and gradient pulse sequences [26, 27]. Therefore, a switchable magnetic field gradient is a requisite for atomic-scale NV-MRI.
In this work, we demonstrate a device tailored for nanoMRI, that provides an on-demand localized magnetic field gradient on the apex of a quartz tip (see Fig. 1a), reaching gradients as high as , as measured by a nearby single NV center. We demonstrate how this tip, deposited with gold, modulates the effective Rabi driving power of the NV center by a factor of , effectively serving as a Rabi power gradient.
2 Results
Current focusing device design
@@ 2 Results
Inspired by the MRFM community [6], the device we present was developed in order to achieve controllable magnetic field gradients localized around a single NV, which acts as a magnetic sensor. To this end, we fabricate a wire by physical vapor deposition (see Methods) that runs along a quartz tip. The abrupt change in the wire’s angle at the apex of the tip results in a gradient in the magnetic field induced by the current running through the wire. The field of such a configuration is estimated by numerical integration according to the Biot-Savart law (see Supplementary Note 1).
@@
The simulation results are presented in Fig. 1(b,c), showing the magnitude of the magnetic field in-plane and out-of-plane relative to the tip. The field strength is weak relative to the zero field splitting, and so this field is not expected to impact the eigenstates of the NV [21]. We considered a working point of 300 nm as a stand-off distance between the tip and the diamond, where the maximum field strength is 0.37 mT (see Supplementary Note 3). The maximum gradient will determine the spatial resolution, calculated to be when running a current of 2 mA. The final resolution of the system depends also on the spectral resolution, or linewidth, of the sensor, for a typical single NV setup at ambient conditions this is around 100 kHz for shallow implantation, corresponding to a dephasing time of of , where ) [28]. Thus, in order to achieve 1 nm spatial resolution for electron spins (with the electron’s gyromagnetic ratio taken to be ) a gradient of is required, similar to the order of magnitude of the calculated gradient.
@@
We found the maximal current in the device is limited due to Joule heating induced by the current, which can cause device failure, i.e., open circuit. In order to estimate the heating of the device, we simulate the heating using COMSOL to solve the heat transfer equations for the device geometry (see Supplementary Note 7). From the perspective of heating, a larger cross-section is beneficial to sustain increasing currents (see Fig. LABEL:fig:temperature), thus a larger apex will prevent device failure due to heating. Nevertheless, our simulations (as shown in Fig. 1) suggest that a smaller apex diameter leads to a stronger gradient. The chosen geometry, therefore, of the device, was set to balance the need for low heating while still achieving a gradient that is beneficial for magnetic resonance imaging. To this end, the tip is designed to have an apex on the order of 1 m and a deposition thickness on the order of 100 nm (see Methods).
The gradient sensing setup
@@
Magnetic field characterization
@@
To map the DC magnetic field of the tip, an optically detected magnetic resonance (ODMR) pulse sequence, sensitive to magnetic fields that are projected along the NV zero-field splitting axis, is used, where current is applied during the spin state manipulation but turned off during the spin readout and initialization phase, see Fig. 2c. We identify a rise/fall time of 600 ns (see Supplementary Note 8 for the MW/tip circuitry). The pulse sequence is alternated with and without current to cancel the effect of slow drifts that would otherwise accumulate and mask small changes in the magnetic field. The tip is scanned relative to the NV and ODMR is measured at each tip location. To avoid damage to the tip, the measurements were performed with a current of 1.54 mA, as measured on the oscilloscope (see Supplementary Note 5 for a discussion on the current amplitude).
@@
The magnetic field measured at a single pixel (tip position) is shown in Fig. 3, with each resonant dip fitted to a Gaussian. The shift in the magnetic field due to the tip is calculated as the difference between the center dip measured with and without the current. The magnetic field is then measured over a lateral range of around the pillar with an NV. Scans performed at varying distances to the diamond’s surface are shown in Fig. 3(a,b). The maximum measured field is 259 , which does not hinder the NV contrast [21]. To compare to the simulation of the expected field (as shown in Fig. 1c), we account for the degrees of freedom in the measurement. The first is the relative angle between the NV axis to the direction of the current, and the other being the lab coordinates along which the tip is scanned. The magnitude and range of the measured field are affected by the angle between the current to the NV, since the ODMR measurement is most sensitive to fields parallel to the NV axis. While the scanning axis defines the axis connecting the maxima and minima, it does not affect the field strength. The simulation, as shown in Fig. 3, projected along the NV axis and rotated by 45∘, is qualitatively similar to the measurement. Several features are visible in both - a sharp peak in the field along with a broad yet shallow dip. As expected, the field strength increases with the distance to the surface. Quantitatively, the range and values of the field in the experiment and the simulation are in good agreement. The discrepancies can be associated with several uncertainties - the discreteness of the measurement, the error in the distance to the surface and simulation assumptions regarding the current flow along the wire. From the comparison, we can then calculate the magnetic field induced by the tip at the sample’s position. This is important when considering the effect of the magnetic field on spins external to the diamond, such as unpaired electrons on the surface and more broadly a sample placed on the diamond, since by knowing the field generated by the gradient at each point, we can deduce what the Larmor frequency of the target spin would be.
@@
To measure the magnetic gradient at higher resolution, we measure ODMR along a line connecting the minima to maxima regions, see Figs. 3c and 4a. This scan was repeated as a function of distance to the diamond surface, as expected the gradient increases as the tip is closer to the diamond surface, see Fig. 4d. Following the law of Biot-Savart (see Supplementary Equation LABEL:eq:_bio-savart), we fit the increasing gradient to an inverse quadratic function. To obtain the maximal gradient the scan should be performed along the optimal x,y curve, where deviations from this curve will result in a sub-optimal gradient as seen in the point closest to the surface (Fig. 4d).
@@
For a current of 1.54 mA through the tip, the maximum measured gradient is , measured when scanning the tip along the axis connecting the minima to maxima regions. The limit on resolution is a combination of the spectral resolution of the setup and the gradient achieved by the tip. With the single NV’s dephasing time () of , we were able to measure ODMR within an error of 100 kHz. The signal from an electron is detected by the NV at the electron resonance frequency at , where and is the magnetic field applied to the electron. Thus, to distinguish electrons in our system, the magnetic field difference required is . While we performed most of the measurements at a relatively low current, a destructive test showed the device can support higher current densities, with a 5-fold improvement in the gradient (see Supplementary Note 5), making it practical for nanometer resolution imaging. In fact, our low-current gradient can already distinguish between spin labels, for example, spaced 4.8 nm in a molecular ruler [30], with a Larmor precession frequency difference between them of 134 kHz, as mentioned above, our gradients can be as high as 3.3 . Additional improvement can be achieved by combining the gradient field with protocols that have demonstrated resolution as good as [31, 32]. Alternatively, the tip itself can be optimized further to improve both geometry and heat load, which may provide a better magnetic gradient.
Tip-induced Rabi power modulation
@@
The geometry used here, i.e., where a metallic tip is placed above the diamond while a metallic microwave waveguide is placed below it, is shown to affect the NV Rabi period dramatically. We observe that by a proper positioning of the metallic tip, the Rabi oscillations’ frequency (, the Rabi driving power) can be increased for a set MW power applied to the waveguide. Fig. 5 shows two sets of measurements demonstrating this effect. First, we scan the distance to the diamond surface for a set lateral position and show a relative, 3.5-fold increase in Rabi power (Fig. 5a). Next, Fig. 5b shows a lateral scan where the tip is set to a distance of 156 nm from the surface, and we plot the ratio between the measured Rabi power, , and the Rabi power when the tip is far-away from the NV, . Here, there are three noticeable regions - (1) the far-field where the Rabi power is unaffected and equal to the baseline Rabi power as measured without a tip; (2) an enhanced region where the Rabi power increases relative to the baseline, and; (3) a third region where the Rabi power is significantly weaker relative to the baseline.
@@
The observed modulation of the Rabi power is due to the metallic tip modulating the magnetostatic near-fields that interact with the NV center. Since these are near-field interactions, the modulation’s strength is very sensitive to the tip’s position. A detailed analytical model, representing the tip geometry using oblate ellipsoids is given in Supplementary Note 9. The contours in Fig. 5b are iso-power lines calculated according to the model, and are in agreement with the experimental data. Our model also qualitatively captures the vertical distance dependence (orange line in Fig. 5a). When simplifying each ellipsoid to a single point dipole, the data can be fitted by a behavior (green curve in Fig. 5a), where is the tip’s vertical distance from the NV and is the lateral one.
@@
3 Discussion
We introduced a magnetic-field gradient on an AFM tip which can be pulsed with rise/fall times shorter than 600 ns. This is achieved using a three-step self-aligning deposition process, effectively creating a focused current at the apex of the tip. Using an NV center as a local magnetic sensor, we demonstrated gradients as strong as , which can be even further increased five-fold with a stronger current. Our proof-of-principle of the gradient enables, already now, nanoscale magnetic resonance imaging of proximate unpaired electrons spins in, e.g., a molecular ruler. Moreover, this metallic tip can also serve as a means to locally modulate the driving field around both the target and sensor (NV) spin, yielding an enhancement of in the spin’s Rabi driving power. The spatial variability of the Rabi power is and of itself a gradient source for future magnetic resonance imaging in nanometer length scales. The geometry of the tip (apex shape and diameter) as well as choice of deposited metal and the substrate material itself can all be further optimized for an overall higher gradient. Specifically, when working at cryogenic temperatures, a superconducting material can support larger current densities [34] (and no Joule heating), the tip’s quartz substrate can be replaced with sapphire or diamond for better heat dissipation [35], and an optimization of the apex geometry while taking into account Joule heating can yield higher magnetic field gradients in proximity to the tip.
The current focusing device offers several advantages, overcoming the limitations of existing nanoMRI techniques. First, the ability to pulse the current allows for different magnetic fields during spin manipulation, readout, and initialization, providing optimal NV contrast as defined by the diamond lattice eigenstates. Second, the magnetic field induced by the tip is small relative to the external field and the zero field splitting, meaning only minor shifts are expected, and so the effective Zeeman energy remains largely unchanged when switching between measurements with and without the field from the tip. The possibility to co-align the tip with an arbitrary NV also holds advantages to sensing when working with single NVs. With the geometry presented here, a sample can be placed on the diamond surface in a statistical manner, and once an optimal NV is identified, the gradient can be applied locally around the specific NV, targeting spins and creating a spectral separation between them in Larmor frequency, in the NV’s immediate vicinity. Additionally, the current source can also reverse the direction of the magnetic field, enabling the application of gradient pulse sequences similar to those used in conventional NMR [26, 27].
The tip combines two critical features for immediate application in nanoMRI, namely control over the temporal and spatial magnetic field. By working at the upper current limit, we can achieve the gradient necessary for nanoscale imaging. The device thus overcomes the limitations of alternative methods, enabling high-contrast spin read-out with an arbitrary choice of the working NV and of sample placement. Going beyond nanoMRI, we envision our scannable tip with a switchable local magnetic field gradient as a useful tool in active and passive characterization of materials, e.g., (anti)ferromagnets, superconductors and more exotic magnetic phases of matter.